The U.S. Environmental Protection Agency set a maximum contaminant level (MCL) for tetrachloroethylene (TCE, C2Cl4) in drinking water of 0.005 ppm (by mass). If the water in your home contains 5 percent of the allowable MCL, what is the molarity of TCE in your home?
1
Expert's answer
2019-09-16T04:03:29-0400

5% of 0.005 ppm is 0.05×0.005= 0.00025 ppm

For water 1ppm≈1mg/L, so we have 0.00025mg/L of TCE.

moles of TCE= 0.00000025 g/168g/mol= 1.5×10^-9 mol


"moles (TCE)= \frac{0.00000025g}{168\frac{g}{mol}}=1.5\times 10^{-9} mol"

Concentration (TCE) ="\frac{1.5\times 10^{-9}}{1L}= 1.5\times 10^{-9} M"